# STABSim: A Parallelized Clifford Simulator with Features Beyond Direct Simulation

### v1

@@ Abstract.
Efficient stabilizer-based computation is a key task in quantum computing as it provides a classically tractable insight into quantum algorithms, which otherwise quickly become impossible at scale on a classical device. Formulations in quantum error correction, hardware verification, hybrid computing, and more, all benefit from the ability of Clifford gates to be accurately evaluated by classical computers. However, current stabilizer simulators have significant limitations in scaling and use-case beyond direct simulation. To address this, we present a GPU-accelerated tableau stabilizer simulator that scales efficiently in direct circuit simulation, and supports a range of stabilizer-based tasks. We show how the simulator is used to quickly calculate Pauli commutation groupings between Pauli strings to reduce the number of circuit runs in quantum chemistry problems. We also present a Clifford+T circuit transpiler based on STABSim, which uses the simulator framework to track relations in non-Clifford rotations as Clifford gates are absorbed into the measurement step of the circuit. The T rotation layers left behind are then further optimized in the simulator structure to create a reduced circuit containing only the minimum non-Clifford rotations, improving the time to simulate and reducing the cost to run on real quantum devices.
@@ 1. Introduction
Quantum computing holds the promise of dramatically accelerating solutions to classically hard problems in areas such as cryptography (Shor, 1999; Gisin et al., 2002), optimization (Farhi et al., 2014; Han and Kim, 2002), quantum chemistry (Georgescu et al., 2014; Kandala et al., 2017), and finance (Rebentrost et al., 2018; Woerner and Egger, 2019), etc. In recent years, hardware demonstrations have shown an increasing number of qubits while reducing error rates. Quantum devices capable of demonstrating of quantum advantage in certain tasks have been reported (Arute et al., 2019; Wu et al., 2021; Gao et al., 2025). However, to gain the full potential of quantum computing, the error of quantum systems needs to be carefully managed and corrected (Nielsen and Chuang, 2010) using quantum error correction (QEC) codes.
QEC encodes dozens of physical qubits into one logical qubit and actively detecting and correcting errors that occur in the quantum systems. QEC lays the foundation for fault-tolerant quantum computing. Early demonstrations of small-scale QEC codes have shown encouraging results (Ryan-Anderson et al., 2021; Krinner et al., 2022; Acharya et al., 2023; Bluvstein et al., 2024; Acharya et al., 2025), but the path toward mid-term and far-term devices featuring thousands of logical qubits demands significant advances in architecture, control, and simulation (Gidney and Ekerå, 2021). In particular, designing and verifying different QEC codes such as surface codes, color codes, and especially understanding the error threshold and the noise performance of logical gate operations, requires efficient large-scale classical simulation tools.
@@
QEC codes primarily consist of Clifford operations, which can be efficiently modeled using the stabilizer formalism (Gottesman, 1998; Nielsen and Chuang, 2010). Unlike general quantum circuit simulators that scale exponentially with qubit count, stabilizer simulators scale polynomially, enabling large-scale QEC design verification. Despite this efficiency, existing stabilizer simulators, including Qiskit (Qiskit Development Team, 2021) and Stim (Gidney, 2021). Qiskit, in addition to its widely used quantum compilation and emulation features, offers a stabilizer simulation backend that can simulate noisy Clifford circuits if all errors in the noise model are also Clifford errors. This has proven beneficial for debugging near-term algorithms and verifying small to mid-sized QEC codes (Goto, 2024; Bravyi et al., 2024a; Cross et al., 2024). Meanwhile, Stim stands out for its meticulous single-threaded CPU optimizations, significantly accelerating large stabilizer circuits through tight memory layouts and specialized bitwise operations (Gidney, 2021). Both Qiskit and Stim highlight the value of specialized stabilizer simulation in quantum computing research, especially for QEC protocols.
However, as we move forward from the Early fault-tolerant era to Megaquop machines (Preskill, 2018; Katabarwa et al., 2024; Preskill, 2025), we expect to expand the simulation probabilities of stabilizer simulators, targeting large-scale multi-logical-qubit code block simulations. However, Stim’s performance often becomes limited by CPU-centric design constraints such as memory bandwidth and the relatively low degree of parallelism available in general-purpose processors. As shown in Figure 1, while Stim performs significantly better than Qiskit, both Qiskit and Stim scale poorly to larger QEC codes.
Large surface-code circuits with repeated rounds of syndrome extraction, fast feedback, and intricate multi-qubit operations—like those in lattice surgery—push the limits of CPU-only approaches. In particular, the sheer volume of stabilizer updates, measurements, and state transformations can quickly overwhelm the available computing resources, resulting in long simulation times and restricting the scale of feasible experiments. In addition, as Stim is optimized towards QEC applications, other applications relying on Clifford and Pauli operations can have difficulties being mapped onto the framework to leverage the power of Stim.
@@
To tackle these challenges, we propose a GPU-accelerated stabilizer simulator that exploits the massive parallelism and high-bandwidth memory of modern Graphics Processing Units (GPUs). Originally designed for graphics workloads, GPUs today provide thousands of concurrent threads and efficient memory subsystems, making them interesting for the large matrix-like operations at the core of QEC simulation. Previous efforts on GPU implementation did not yield performance improvement (Gidney, 2021), citing the low arithmetic intensity of XOR operations at the core of stabilizer simulation. CPUs certainty outperform GPUs for small and sequential arithmetic by design. However, dedicated GPUs have their own bitwise and synchronization capabilities that can safely enable many operations at once, even in the time-order sensitive context of quantum circuit simulation. Furthermore, comparisons and longer arithmetic operations appear frequently during measurement operations, which are the bottleneck of QEC simulation times.
Our approach encodes stabilizer states directly on the GPU, orchestrating operations so that they can be applied in parallel, thereby significantly reducing simulation time and time scaling for larger circuits. This design enables studies of mid-term and far-term QEC protocols, including high-distance surface codes and complex lattice surgery which scales impractically on lightly-threaded CPUs. We still provide lightweight CPU version which performs within the realm of other modern stabilizer simulators. Both versions are flexible with fast stabilizer and Pauli-based computing features beyond direct simulation. As we will show, efficient stabilizer simulation has useful applications for fields beyond QEC, including chemistry simulation and circuit optimization.
This work will detail the following contributions:
@@
    The first quantum stabilizer simulator to show performance gains via GPU in Clifford+Measurement circuits.
@@
    Improved scaling among stabilizer simulators as qubit size increases, lowering the time to evaluate a full surface code simulation and measurement after $\sim 2000$ qubits.
@@
    Integrated Pauli-Based Computing (PBC) capabilities and a direct quantum chemistry use-case shown for VQE.
@@
    A novel Clifford+T to PBC transpiler using the stabilizer simulator, showing how stabilizer simulation can be leveraged to improve non-Clifford circuits by reducing T-gate state count by 35-70%. This transpiler is the first to our knowledge to automatically convert any static Clifford+T circuit to a PBC circuit using bitwise methods, and the first to perform T-state optimization.
@@ 2.1. Clifford Gates
At the core of most quantum computing circuits are Pauli operators, which consist of tensor products of Pauli operators $X$, $Y$, $Z$, and identity operator $I$. With the combination of phase factors {±$\{\pm 1,\pm i\}$, Pauli operators form Pauli groups across $n$ qubits.
The Clifford group comprises all unitary operators that map Pauli operators to other Pauli operators under conjugation. In practice, this group is generated by a small set of gates, e.g., single-qubit Hadamard gates $H$, phase gate $S$, and $CX$ gates, represented in computational basis as
(1) $H=\frac{1}{\sqrt{2}}\begin{pmatrix}1&1\\ 1&-1\end{pmatrix},\,S=\begin{pmatrix}1&0\\ 0&i\end{pmatrix},\,\mathrm{CX}=\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&0&1\\ 0&0&1&0\end{pmatrix}.$
Proved by the Gottesman-Knill theorem, quantum circuits restricted to Clifford gates and Pauli measurements can be simulated on a classical computer in polynomial time with respect to the circuit size and number of qubits (Aaronson and Gottesman, 2004; Gottesman, 1998; Nielsen and Chuang, 2010). Luckily for us, QEC and many other applications are formulated in a way that Clifford gates are sufficient.
@@ 2.2. The Stabilizer formalism
A stabilizer of a quantum state $\lvert\psi\rangle$ is an operator $\hat{A}$ such that
(2) $\hat{A}\ket{\psi}=+1\ket{\psi}$
Although stabilizers can be chosen from broader operator sets, Pauli stabilizers (i.e., tensor products of the single-qubit operators I, X, Y, Z) are especially popular due to their simple algebraic structure. A stabilizer group is a commuting (Abelian) subgroup of the $n$-qubit Pauli group whose elements all square to the identity and do not include the negative identity operator $-I$. A state stabilized by a stabilizer group is a stabilizer state. Stabilizer states can be equivalently described by the stabilizers in the stabilizer group, especially the group generators.
The effect of Clifford group operations on stabilizer states can be simulated by tracking the transformation of the corresponding stabilizer group. Pauli measurement can also be tracked using the following rules (Litinski, 2019),
@@
    if the operator is not in the stabilizer group $\mathcal{S}$, the measurement outcome can be either ±$\pm 1$. The state after the measurement is stabilized by a new group that contains the new measured operator ±$\pm g$ depending on the measurement outcome and all other group generators that commute with $g$ in the original stabilizer group $\mathcal{S}$.
A stabilizer $\mathcal{S}$ represents one possible tensor product of Pauli operators that stabilize a state by equation 2. However, a pure $n$-qubit stabilizer state is stabilized by $n$ commuting and independent Pauli stabilizers. All of these stabilizers can be listed in a table, where rows are the independent stabilizers, and columns are the qubit index.
@@
In order to efficiently track the transformation of Pauli operators, Pauli operators are frequently represented by binary vectors. Each Pauli operator $\sigma$ is represented by two bits, an $x$ portion and a $z$ portion. Pauli $X$ is represented by $(1,0)$, $Z\mapsto(0,1)$, $Y\mapsto(1,1)$, and $I\mapsto(0,0)$. The commutation relation of two Pauli operators represented as $(x_{1},z_{1})$ and $(x_{2},z_{2})$ can also be easily computed using
(4) $(x_{1}\cdot z_{2})\oplus(x_{2}\cdot z_{1}),$
where $\oplus$ is an xor, and a result of 0 means the two Pauli operators commute.
@@ 2.3. The CHP Formalism
Aaronson and Gottesman proposed an optimized version of the tableau, specifically for optimizing the measurement operations, known as the CNOT-Hadamard-Phase (CHP) formalism (Aaronson and Gottesman, 2004). For a $n$-qubit state, instead of a single tableau that only tracks the $n$ stabilizer group generators, in CHP formalism, an extra tableau is to added to track the $n$ destabilizers. Specifically, the tableau that are tracked in the CHP formalism can be represented as
(5) $\left(\begin{array}[]{ccc|ccc|c}x_{11}&\cdots&x_{1n}&z_{11}&\cdots&z_{1n}&r_{1}\\ \vdots&\ddots&\vdots&\vdots&\ddots&\vdots&\vdots\\ x_{n1}&\cdots&x_{nn}&z_{n1}&\cdots&z_{nn}&r_{n}\\ \hline\cr x_{(n+1)1}&\cdots&x_{(n+1)n}&z_{(n+1)1}&\cdots&z_{(n+1)n}&r_{n+1}\\ \vdots&\ddots&\vdots&\vdots&\ddots&\vdots&\vdots\\ x_{(2n)1}&\cdots&x_{(2n)n}&z_{(2n)1}&\cdots&z_{(2n)n}&r_{2n}\end{array}\right)$
where each row represents a stabilizer or destabilizer. The first $n$ rows correspond to the stabilizers, while the last $n$ rows are the destabilizers. The columns $r$ represent the sign bits of each Pauli operator. By tracking the destabilizers, it removes the need to do Gaussian elimination on the stabilizer tableau when a Pauli measurement is performed, reducing computational complexity from $O(n^{3})$ to $O(n^{2})$ (Aaronson and Gottesman, 2004). The tableau requires $4n^{2}+n$ bits to store all necessary stabilizers and destabilizers, including a phase bit for each, for CHP operations. Thankfully, this means that a full tableau representation of systems with tens of thousands of qubits still only requires space on the order of MB. For comparison, a state vector or density matrix representation of the state will grow on the order $O(2^{n})$. These simulation methods become infeasible around roughly 50 qubits (Chen et al., 2025).
@@ 2.4. Operations on the CHP Tableau
Once the data structure of the tableau is established, we want to perform minimal bitwise operations to simulate the evolution of the tableau under Clifford operations. Clifford gates transform Pauli operators to Pauli operators under conjugation. For example, $HXH^{\dagger}=Z$. From this, some basic bit transforms on the tableau can be calculated that represent the action of Clifford gates on the stabilizers of a system. Most often, all stabilizers in the tableau will receive the same transformation gate-by-gate, meaning Clifford gate operations can be similar to matrix operations. However, the most time consuming operations, measurements, require more logic.
Gate Binary Tableau Operation
@@
Measure ($M$) $\begin{cases}rowsum(i,p)\{\forall i\mid x_{i,q}=1\}&\text{if }random\\ rowsum(temp,i){\forall i>n}&\text{if }deterministic\\ \end{cases}$
@@
Table 1. Clifford gates and their bit operations on the stabilizer tableau. Hadamard, phase and CNOT operations occur on every row of the tableau, at the column that the gate targets. c and t represent the control and target bit of a CNOT operation.

### v2

@@ Abstract.
The quantum stabilizer formalism became foundational for understanding error correction soon after the realization of the first useful quantum error correction codes. Stabilizers provide a way to describe sets of quantum states which are valid codewords within a quantum error correction (QEC) scheme. Existing stabilizer simulators are single threaded applications used to sample larger codes than is possible with other methods. However, there is an outstanding gap in the scaling and accuracy of current simulators for QEC as quantum computing exceeds hundreds of qubits, along with an under-utilization of the capabilities of highly-efficient stabilizer simulation across other quantum domains. In this work, we present the first GPU-accelerated tableau stabilizer simulator to scale better than CPU methods in QEC workloads, by trivializing Clifford gates and exploiting the large parallelism of dedicated GPUs with CUDA warp-level primitives to quickly overcome costly measurement gates. We then implement a new error model that captures non-unitarity in T1/T2 error channels much faster and with exact accuracy for most physical qubits, demonstrate a chemistry use case, and present a new Clifford+T to Pauli-Based Computing (PBC) transpilation optimization through our simulator.
@@ 1. Introduction
Quantum computing holds the promise of dramatically accelerating solutions to classically hard problems in areas such as cryptography (Shor, 1999; Gisin et al., 2002), optimization (Farhi et al., 2014; Han and Kim, 2002), quantum chemistry (Georgescu et al., 2014; Kandala et al., 2017), etc. In recent years, hardware demonstrations have shown an increasing number of qubits while reducing error rates. Quantum devices capable of demonstrating of quantum advantage in certain tasks have been reported (Arute et al., 2019; Wu et al., 2021; Gao et al., 2025). However, to gain the full potential of quantum computing, the error of quantum systems needs to be carefully managed and corrected (Nielsen and Chuang, 2010) using QEC codes.
QEC encodes many physical qubits into one logical qubit and actively detects and corrects errors that occur in the quantum systems. QEC lays the foundation for fault-tolerant quantum computing. Early demonstrations of small-scale QEC codes have shown encouraging results (Ryan-Anderson et al., 2021; Krinner et al., 2022; Acharya et al., 2023; Bluvstein et al., 2024; Acharya et al., 2025), but the path toward mid-term and far-term devices featuring thousands of logical qubits demands significant advances in architecture, control, and simulation (Gidney and Ekerå, 2021). In particular, designing different QEC schemes while understanding the error threshold and noise performance of logical gate operations, requires very efficient classical simulation tools.
@@
QEC codes primarily consist of Clifford operations and measurements, which can be efficiently modeled using the stabilizer formalism (Gottesman, 1998; Nielsen and Chuang, 2010). Unlike general quantum circuit simulators that scale exponentially with qubit count, stabilizer simulators scale polynomially, enabling large-scale QEC design verification. Qiskit, in addition to its widely used quantum compilation and emulation features, offers a stabilizer simulation backend that can simulate noisy Clifford circuits if all errors in the noise model are also Clifford errors. This has proven beneficial for debugging near-term algorithms and verifying small to mid-sized QEC codes (Goto, 2024; Bravyi et al., 2024a; Cross et al., 2024). Meanwhile, Stim stands out for its meticulous single-threaded CPU optimizations, significantly accelerating large stabilizer circuits through tight memory layouts and specialized bitwise operations (Gidney, 2021). Both Qiskit and Stim highlight the value of specialized stabilizer simulation in quantum computing research, especially for QEC protocols.
However, as we move forward from the early fault-tolerant era to Megaquop machines (Preskill, 2018; Katabarwa et al., 2024; Preskill, 2025), we expect to expand the simulation capabilities of stabilizer simulators, targeting large-scale multi-logical-qubit code block simulations, with individual patch and multipatch code distances from tens to hundreds. In this regime, the polynomial overhead becomes a challenge for state-of-the-art simulators. Large QEC circuits with repeated rounds of syndrome extraction, fast feedback, and many qubit operations push the limits of CPU-only approaches. Operations like lattice surgery are key for fault tolerance, and have modern tools such as TQEC to enable the construction of enormous fault-tolerant pipelines (noa, 2025).
However, CPU-centric design constraints such as memory bandwidth and low parallelism, place a practical time constraint on actually characterizing errors in these huge architectural constructions. As shown in Figure 1, while Stim is significantly better than Qiskit, both Qiskit and Stim scale poorly to larger QEC codes. Furthermore, to take advantage of Stim’s efficiency via Pauli frames, physical qubit errors must be approximated using Pauli operations, which is an assumption that quickly diverges from reality when considering common non-unitary errors such as qubit relaxation.
Particularly, the volume of stabilizer updates across the stabilizer tableau and scaling faster per-qubit measurement costs quickly overwhelm sequential CPU pipelines, resulting in long simulation times and restricting the scale of practical investigation. In addition, as Stim is highly optimized for shot-based QEC evaluation, other applications relying on Clifford and Pauli operations are difficult to integrate.
@@
To tackle these challenges, we propose a GPU-accelerated stabilizer simulator that exploits the massive parallelism and high-bandwidth memory of modern Graphics Processing Units (GPUs). Originally designed for graphics workloads, GPUs today provide thousands of concurrent threads and efficient memory subsystems, making them an interesting platform for the large matrix-like operations at the core of QEC simulation. Previous efforts on GPU implementation either did not yield performance improvement (Gidney, 2021), or could not perform measurement, which is the key operation to QEC (Osama et al., 2025). Stim cited the low arithmetic intensity of bit operations at the core of stabilizer simulation as the reason for GPU under-performance. For measurement, the strict time-order sensitivity and cost of branching posed a barrier that has not been addressed until this work.
Our approach encodes stabilizer states directly on the GPU, orchestrating operations to be applied in parallel with minimal synchronization. Where time-order synchronization is absolutely necessary, extremely efficient CUDA warp-level primitives and optimized block-wide cooperations are employed, allowing us to rapidly clear synchronization and branching bottlenecks. These techniques greatly reduce simulation time as the qubit number and circuit depth increase. This design enables studies of large-scale QEC protocols, including high-distance QEC codes and complex lattice surgery operations which scale impractically on lightly-threaded CPUs. We still provide a CPU version comparable with existing stabilizer simulators as a flexible platform for validating new stabilizer-based algorithms. Both versions support and have been used to develop Pauli-based computing features well beyond direct simulation. As we will show, efficient stabilizer simulation is necessary for large sale QEC, but also has surprising applications for better qubit noise modeling, chemistry simulation, and quantum circuit optimization. This work thus detail four core contributions:
@@
    The first quantum stabilizer simulator to show a performance gain via GPU in Clifford+Measurement simulation. Within a hundreds of qubits, STABSim exceeds state-of-the-art simulators in QEC experiments thanks to highly-efficient CUDA warp primitives combined with careful synchronization for accuracy.
@@
    Implementation of a new qubit noise model that can be sampled with much more efficiency and accuracy than existing simulators. When $T1$ coherence equals or exceeds $T2$, where most physical qubits reside, our implementation only requires tableau sampling with constant overhead, but has exact accuracy to the physical noise channel. Even when $T2$ is greater, our implementation still requires much fewer tableau samples than other exact methods.
@@
    A quantum chemistry use-case for VQE using fast bitwise Pauli string methods available on STABSim.
@@
    A novel Clifford+T to PBC transpiler using the stabilizer framework. We detail how stabilizer simulation can improve universal circuit architectures by optimizing T-gate count, and show how our method greatly improves the speed of transpilation versus other frameworks.
@@ 2.1. Clifford Gates
At the core of most quantum computing circuits are Pauli operators, which consist of tensor products of Pauli operators $X$, $Y$, $Z$, and identity operator $I$. Pauli operators, combined with phase factors $\{\pm 1,\pm i\}$, form Pauli groups across $n$ qubits.
The Clifford group comprises all unitary operators that map Pauli operators to other Pauli operators under conjugation. Proved by the Gottesman-Knill theorem, quantum circuits restricted to Clifford gates and Pauli measurements can be simulated on a classical computer in polynomial time with respect to the circuit size and number of qubits (Aaronson and Gottesman, 2004; Gottesman, 1998; Nielsen and Chuang, 2010). Luckily for us, QEC and many other applications are formulated in a way that Clifford gates are sufficient.
@@ 2.2. The Stabilizer formalism
[…]
The effect of Clifford group operations on stabilizer states can be simulated by tracking the transformation of the corresponding stabilizer group. Pauli measurement can also be tracked using the following rules (Litinski, 2019):
@@
    If the operator is not in the stabilizer group $\mathcal{S}$, the measurement outcome can be either $\pm 1$. The state after the measurement is stabilized by a new group that contains the new measured operator $\pm g$ depending on the measurement outcome and all other group generators that commute with $g$ in the original stabilizer group $\mathcal{S}$.
A stabilizer $\mathcal{S}$ represents one possible tensor product of Pauli operators that stabilize a state. However, a pure $n$-qubit stabilizer state is stabilized by $n$ commuting and independent Pauli stabilizers. All of these stabilizers can be listed in a table, where rows are the independent stabilizers, and columns are the qubit index.
@@
In order to efficiently track the transformation of Pauli operators, Pauli operators are frequently represented by binary vectors. Each Pauli operator $\sigma$ is represented by two bits, an $x$ portion and a $z$ portion. Pauli $X$ is represented by $(1,0)$, $Z\mapsto(0,1)$, $Y\mapsto(1,1)$, and $I\mapsto(0,0)$. The commutation relation of two Pauli operators represented as $(x_{1},z_{1})$ and $(x_{2},z_{2})$ can also be easily computed using $(x_{1}\cdot z_{2})\oplus(x_{2}\cdot z_{1})$, where $\oplus$ is an xor, and a result of 0 means the two Pauli operators commute.
@@ 2.3. The CHP Formalism
Aaronson and Gottesman proposed an optimized version of the stabilizer tableau for optimizing measurements, known as the CNOT-Hadamard-Phase (CHP) formalism (Aaronson and Gottesman, 2004). For an $n$-qubit state, instead of a single tableau that only tracks the $n$ stabilizer group generators, in CHP formalism, an extra tableau is added to track the $n$ destabilizers. Specifically, the tableau that are tracked in the CHP formalism can be represented as
$\left(\begin{array}[]{ccc|ccc|c}x_{11}&\cdots&x_{1n}&z_{11}&\cdots&z_{1n}&r_{1}\\ \vdots&\ddots&\vdots&\vdots&\ddots&\vdots&\vdots\\ x_{n1}&\cdots&x_{nn}&z_{n1}&\cdots&z_{nn}&r_{n}\\ \hline\cr x_{(n+1)1}&\cdots&x_{(n+1)n}&z_{(n+1)1}&\cdots&z_{(n+1)n}&r_{n+1}\\ \vdots&\ddots&\vdots&\vdots&\ddots&\vdots&\vdots\\ x_{(2n)1}&\cdots&x_{(2n)n}&z_{(2n)1}&\cdots&z_{(2n)n}&r_{2n}\end{array}\right)$
where each row represents a stabilizer or destabilizer. The first $n$ rows correspond to the destabilizers, which eliminate the need for $O(n^{3})$ Gaussian elimination (Aaronson and Gottesman, 2004), while the last $n$ rows are the stabilizers. The columns $r$ represent the sign bits of each Pauli string (stabilizer). The tableau requires $4n^{2}+2n$ bits to store all necessary stabilizers and destabilizers, including a phase bit for each, for CHP operations. Thankfully, this means that a full tableau representation of systems with tens of thousands of qubits still only requires space on the order of MB. For comparison, a state vector or density matrix representation of the state will grow on the order $O(2^{n})$ and $O(4^{n})$ respectively, becoming infeasible with a few dozen qubits (Chen et al., 2025).
@@ 2.4. Operations on the CHP Tableau
Once the data structure of the tableau is established, we want to perform minimal bitwise operations to simulate the evolution of the tableau under Clifford operations. Clifford gates transform Pauli operators to Pauli operators under conjugation. For example, $HXH^{\dagger}=Z$. Some basic bit transforms on the tableau can be calculated that represent the action of Clifford gates on the stabilizers of a system. Most often, all stabilizers in the tableau will receive the same transformation gate-by-gate, meaning Clifford gate operations can be similar to matrix operations. However, the most time-consuming operations, i.e., measurements and resets, require much more expensive logic.
Gate Binary Tableau Operation (Aaronson and Gottesman, 2004)
@@
[…]
@@
Table 1. Clifford gates and their bit operations on the stabilizer tableau. Hadamard, phase and CNOT operations occur on every row of the tableau, at the column that the gate targets. c and t represent the control and target bit of a CNOT operation. The original rowsum algorithm is detailed in (Aaronson and Gottesman, 2004) and also provided in Appendix A.2 for convenience.